three babies are born in a family.find the probability
1...a family has 3 male babies given that all the 3 babies have same gender
2....a family has 2 male babies when it is known that first baby is male

Answer by stanbon(75887)

three babies are born in a family.find the probability
1...a family has 3 male babies given that all the 3 babies have same gender
Ans: 1/2
-------------------
2....a family has 2 male babies when it is known that first baby is male
Ans: P(one male in 2 births) = 2C1(1/2)^1*(1/2)^1 = 2*1/4 = 1/2
====================
Cheers,
Stan H.